OpenStudy (r_josh10):

A truck travels 239 miles in 4.2 hours. What is his average speed? A 1003.8 mi/hr B 113.8 mi/hr C 56.9 mi/hr D 0.04 mi/hr

OpenStudy (r_josh10):

@dan815

OpenStudy (r_josh10):

@paki

OpenStudy (r_josh10):

@texaschic101

OpenStudy (anonymous):

wouldn't you multiply 239 * 4.2 to get the answer?

OpenStudy (r_josh10):

I think the formula is S=D/T

OpenStudy (mrnood):

@RavenTeenTitans NO you divide miles/hours to get mph

OpenStudy (r_josh10):

is the formula right tho? @MrNood

OpenStudy (anonymous):

Oh right, the formula. I'm becoming more and more forgetful each day.

OpenStudy (r_josh10):

np @RavenTeenTitans

OpenStudy (r_josh10):

Is it C? Since S=D/T, S=239/4.2?

OpenStudy (mrnood):

well the formula IS speed = distance/time but this is usually written v=s/t using s for distance, t for time and v for velocity (or speed in this case)

OpenStudy (xmissalycatx):

You're going to divide the two numbers because your formula would be \[v=\frac{ s }{ t }\] so it would look like.. \[v=\frac{ 239 }{ 4.2 }\]

OpenStudy (r_josh10):

so is C right?

OpenStudy (xmissalycatx):

You're correct. :)

OpenStudy (r_josh10):

Thanks! @xMissAlyCatx

OpenStudy (mrnood):

if you understand theconcept and are confident in your calculation the just go with it. There is not always someone to confirm tyour work in 'real life'

OpenStudy (xmissalycatx):

You're welcome @r_josh10 and @MrNood is right. If you really are confident in your answer then go with it. But we will be here to assist you if you need it! :)